OpenStudy (anonymous):

A sphere is increasing in volume at the rate 3pi cm^3/s. At what rate is its radius changing when the radius is 1/2 cm?

OpenStudy (anonymous):

V=4/3 pi r^3 dV/dt=d(4/3pi r^3)/dt

OpenStudy (anonymous):

Now, use chain rule to find dr/dt

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

dr/dt = 1/(4pi r^2) * (dV/dt)

OpenStudy (anonymous):

Substitute 1/2 in for r?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and dV/dt=3pi cm^3/s

OpenStudy (anonymous):

dr/dt = 1/(4pi*((1/2)^2))*3pi cm^3/s = 3 cm^3/s Thank you!! :)

OpenStudy (anonymous):

welcome